OpenStudy (anonymous):

Write the equation of the line which passes through (2, -3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

the slope of y = 4x + 7 is 4, so the slope of a line perpendicular would be -1/4 using point-slope form, y-y1 = m(x-x1) y +3 = -1/4(x-2) multiply both sides by -4 to get rid of the fraction -4y - 12 = x - 2 rearrange to get x + 4y = -10

OpenStudy (anonymous):

Thank You! :)